A polarised & an analyser are oriented so that the maximum light is transmitted. What is the fraction of maximum light transmitted when analyser is rotated through (1) 30 degree (2) 60 degree ?

Grade:12

1 Answers

Arun
6 years ago
Dear Harshita
since
I=I0/2(when light pas through polaraiser)
and
I'=I cos2theta
so
(i)
finally we have
I'=(I/2)cos²30°   
=(I/2)(3/4)
=(3/8)
ii)
I = (I/2) cos² 60°
= (I/2) (1/4)
= 1/8
Regards
Arun